### v1

# JWST detection of a carbon dioxide dominated gas coma surrounding interstellar object 3I/ATLAS

Abstract
3I/ATLAS is the third confirmed interstellar object to visit our Solar System, and only the second to display a clear coma. Infrared spectroscopy with the James Webb Space Telescope (JWST) provides the opportunity to measure its coma composition and determine the primary activity drivers. We report the first results from our JWST NIRSpec campaign for 3I/ATLAS, at an inbound heliocentric distance of au. The spectral images (spanning – m) reveal a CO2 dominated coma, with enhanced outgassing in the sunward direction, and the presence of H2O, CO, OCS, water ice and dust. The coma CO2/H2O mixing ratio of is among the highest ever observed in a comet, and is above the trend as a function of for long-period and Jupiter-family comets (excluding the outlier C/2016 R2). Our observations are compatible with an intrinsically CO2-rich nucleus, which may indicate that 3I/ATLAS contains ices exposed to higher levels of radiation than Solar System comets, or that it formed close to the CO2 ice line in its parent protoplanetary disk. A low coma H2O gas abundance may also be implied, for example, due to inhibited heat penetration into the nucleus, which could suppress the H2O sublimation rate relative to CO2 and CO.
1 Introduction
Comets and planetesimals are theorized to form in large numbers during the accretion of planetary systems. Many of these small bodies, composed of ice, rock, and dust, are subsequently expected to be ejected into interstellar space through gravitational encounters with larger (planetary or stellar) bodies (L. Dones et al., 2004; H. F. Levison et al., 2010; S. N. Raymond et al., 2018; S. Pfalzner & K. Vincke, 2020; X.-L. Zheng & J.-L. Zhou, 2025). The apparitions of the first confirmed interstellar objects (ISOs) 1I/‘Oumuamua in October 2017 and 2I/Borisov in August 2019 provided confirmation of this theory, and offered an unprecedented opportunity to study the nature of matter delivered to our Solar System from a distant planetary system, thus spawning a new field of planetary science.
Due to its faintness and short observing window, spectroscopic characterization of volatiles from 1I/‘Oumuamua proved elusive, so the composition of this object remains highly uncertain (‘Oumuamua ISSI Team et al., 2019). On the other hand, for the intrinsically brighter and more active 2I/Borisov, ultraviolet, optical, and submillimeter observations (e.g. M. A. Cordiner et al., 2020; D. Bodewits et al., 2020; P. Guzik & M. Drahus, 2021; C. Opitom et al., 2021; Z. Xing et al., 2020) provided tantalizing glimpses of its coma composition, and revealed an object that was similar in many ways to the well-studied comets from our own Solar System, but with an unusually strong enrichment in carbon monoxide (CO) (M. A. Cordiner et al., 2020; D. Bodewits et al., 2020). Considering the difficulty of studying the ices in the midplanes of protoplanetary disks and planetary systems elsewhere in our Galaxy, continued spectroscopic observations of interstellar objects have the potential to reveal crucial details on the physics and chemistry of planet formation in planetary systems other than our own.
The discovery of a third interstellar object (3I/ATLAS) was announced on 2025 July 1 by the Asteroid Terrestrial-impact Last Alert System (ATLAS; J. L. Tonry et al., 2018). Based on its inbound orbital eccentricity () and heliocentric radial velocity projected to infinity ( km s-1) (D. Z. Seligman et al., 2025), 3I/ATLAS has been confirmed to be on a gravitationally unbound, hyperbolic interstellar trajectory. Dynamical modeling of a population of Galactic interstellar objects (ISOs) shows that the high velocity of 3I/ATLAS is consistent with a relatively large dynamical age of 3–11 billion years (M. J. Hopkins et al., 2025; A. G. Taylor & D. Z. Seligman, 2025). This age, coupled with its trajectory, implies 3I/ATLAS could have originated from a relatively old, low-metallicity stellar system, plausibly from the kinematically hot, “thick disk” population of the Milky Way. Chemical differences between the volatile content of 3I/ATLAS and our Solar System’s comets may therefore be expected.
Similar to 2I/Borisov, 3I/ATLAS has been shown to display clear cometary activity (D. Jewitt & J. Luu, 2025; M. R. Alarcon et al., 2025; M. Minev et al., 2025). Early spectroscopic and photometric observations have revealed a compact nucleus (effective radius km), a bright, dusty coma, with a dust mass-loss rate of 12–120 kg s-1 (D. Jewitt et al., 2025a), and a red spectral slope, with possible water ice absorption at 2.0 µm (T. Kareta et al., in press; B. Yang et al., 2025). The Neil Gehrels-Swift Observatory detected ultraviolet emission from the gas-phase OH radical (Z. Xing et al., 2025), and assuming H2O to be the photolysis parent, a water production rate of s-1 was derived at a heliocentric distance of au.
Here, we present the first infrared spectroscopic observations from our campaign to observe 3I/ATLAS using the James Webb Space Telescope (JWST). This article focuses on analyzing the rovibrational fluorescence emission from H2O, CO2, and CO, which are the main drivers of coma activity in typical Solar System comets. Molecular production rates and mixing ratios are derived, enabling the first chemical characterization of the object’s volatile gas inventory.
2 Observations
Observations of 3I/ATLAS were performed using JWST (J. P. Gardner et al., 2023) on UT 2025-08-06 between 11:02–11:20, using the NIRSpec integral field unit (IFU), as part of program ID 5094. The IFU (T. Böker et al., 2022) divides its field of view into 30 parallel slices, each measuring . The light from each slice is then passed through a dispersing element (in this case, a fluorite prism), to produce 30 contiguous spectra on the NIRSpec detector. Subsequent image processing divides each IFU slice into 30 of spaxels, resulting in a spectral cube, covering the entire field of view. For reference, the FWHM of the JWST point-spread function is at 3 m. Each spectrum covers –5.3 m, with a resolving power that varies from at 1.2 m to at 5.3 m.
3I/ATLAS was acquired in the IFU and tracked using JPL Horizons ephemeris solution , which was constructed based on astrometry supplied by a large number of optical observers. The total on-target exposure duration was 640 s, divided across four dither positions, each spatially separated (in the approximate shape of a square) with offsets of from the (central) target position. The data were reduced using the JWST Calibration Pipeline software version v1.19.1 (H. Bushouse et al., 2025) using the JWST Calibration Reference Data System context file 1413. The four dithers were shifted and combined in the rest frame of the comet during image processing, thus allowing detector artifacts and cosmic rays to be identified and removed. A similar set of four exposures of the sky background was obtained, offset by along the horizontal axis of the IFU aperture. This allows contamination from background infrared sources, zodiacal light, and the telescope to be identified and subtracted. Data cubes with and without background subtraction were produced. The latter allows for the analysis of faint gas emission bands without the contribution of additional noise from a background subtraction. For each spectral data cube, the pipeline produces uncertainty and data quality maps, which were used during our analysis, in particular, for the derivation of formal () error estimates. After combining the four dithered observations, the absolute calibration accuracy is expected to be 3%.111https://jwst-docs.stsci.edu/jwst-calibration-status/nirspec-calibration-status/nirspec-ifu-calibration-status
3 Results
The observed flux was integrated over the entire IFU field of view to produce the spectrum shown in Figure 1. Prominent features include a broad maximum at around 1.2 m, due to scattered sunlight from coma dust grains, and a strong, narrow (double) peak at 4.3 m, which is assigned to the main () rovibrational emission band of gas-phase CO2. Weaker gas emission bands from H2O (), CO () and 13CO2 () are also present, along with broad absorption features centered around 3.0 m and 4.5 µm, attributed to the OH stretching mode and lattice vibrations, respectively, of H2O ice in the coma — likely in the form of small ( micron-sized) icy grains (M. H. Moore & R. L. Hudson, 1992; G. Leto & G. A. Baratta, 2003; R. M. Mastrapa et al., 2009).
To isolate the CO2, H2O, and CO gas emission features, spectral data were extracted for each IFU pixel within the vicinity of each emission band, and a polynomial continuum fit was performed, excluding the spectral region directly inside each molecular emission band (see Figure 4). For CO2 and CO, a 3rd-order polynomial was used, whereas for H2O a 5th-order polynomial was used, to better fit the wing of the 3.0 m ice band (see Section 4). After continuum subtraction, the IFU spectra were integrated across the detected full emission width of each feature, then plotted as maps in Figure 2. The spatially averaged, continuum-subtracted emission band profile for each molecule is shown in the upper-right inset of their respective panels. The m scattered light image (integrated between 0.8–1.4 m) is also shown. The continuum brightness at 0.75 and 1.25 µm is 1013 and 1244 Jy inside a -diameter circular aperture, corresponding to cometary quantities (M. F. A’Hearn et al., 1984) of 392 and 492 cm, respectively (quoted without a correction for phase darkening).
The gas and dust maps for 3I/ATLAS show a well-defined peak, offset East from the center of the IFU by . As shown by a later ephemeris reconstruction, about of the offset can be explained by a error on the predicted ephemeris, while the remaining offset remains under investigation. The 1.2 m scattered light and CO2 emission maps reveal an extended coma of dust and gas that spans the full extent of the NIRSpec IFU. The H2O and CO emission is weaker, and therefore noisier, but nevertheless confirms the presence of a spatially extended molecular coma. While the scattered light shows a clear asymmetry along the Sun-comet axis — enhanced in the direction of the sky-projected comet-Sun and velocity vectors (see also D. Jewitt et al. 2025b; C. O. Chandler et al. 2025) — the gas distributions appear relatively more symmetrical.
To further investigate the coma structure, a enhanced version of the dust map (where is the sky-projected nucleocentric distance), is shown in Figure 3, and similarly for the gas maps in Appendix A (Figure 7). In these enhanced maps, the dominant coma spatial feature: dilution of the observed column densities due to quasi-spherical expansion, has been divided out. The -enhanced 1.2 m map reveals a strong, plume-like feature emanating from the pseudo-nucleus in the approximate direction of the Sun (slightly north of west), with an additional, weaker feature to the north-east. Since the dust in relatively faint comets such as 3I/ATLAS is optically thin at this wavelength, the shape of this feature is interpreted as resulting from an enhanced coma dust density in the sunward direction, potentially from the fragmentation of dust grains increasing the scattering cross section with distance from the nucleus. The gas maps on the other hand, show more subtle asymmetries. CO2 exhibits an azimuthal minimum towards the north, that we attribute to a combination of weak coma sub-structure and optical depth effects, since CO2 becomes optically thick close to the nucleus.
Taking the ratio of continuum-subtracted fluxes within a circular aperture placed immediately adjacent to the sunward () and antisunward () sides of the brightest pixel (the assumed location of the nucleus) gives for the (1.2 m) dust, 1.07 for CO2, 1.04 for CO, and 1.20 for H2O (with uncertainties of on all measurements). Our observations thus reveal a heterogeneous coma morphology consistent with different outgassing patterns for the different species. Such heterogeneity can be explained by various factors, such as the different molecular sublimation temperatures, nucleus release and coma acceleration mechanisms, as well as differing inertial/fluid-dynamical properties for the gas and dust. The degree of asymmetry of the observed gases is likely related to their sublimation temperatures (), with , congruent with (M. Womack et al., 2017). This implies that the sublimation of CO2 and CO is more fully activated than H2O. Full interpretation of the observed outgassing morphologies, including the surprisingly strong sunward dust enhancement, will require detailed physical modeling. Nevertheless, our data are consistent with the origin of this dust feature being influenced by enhanced gas sublimation (and therefore, outgassing) rates associated with the higher temperature on the dayside of the nucleus.
The 3.0 µm H2O ice band depth is 78.30.2% at 2.9–3.1 µm, with respect to the continuum at 2.5 and 3.8 µm. The band remains strong across the field of view, varying by no more than 10%. Water ice also has broad near-infrared absorption features at 1.5, 2.0, 3.0, and 4.5 µm. The 4.5 µm band is difficult to measure in our data due to blending with CO2 and CO gas emission and likely thermal continuum, the 2.0 µm band appears very weak, and there is no discernible 1.5 µm band. Furthermore, the 2.0 µm spectral region in our combined IFU dataset is affected by several bad pixels. In two dither positions with clean spectra in this region, the 2.0 µm band depth is (2.00.1)% and (1.60.1)%. This was estimated by normalizing the reflectance spectrum within a 04 radius aperture with a linear fit between 1.75 and 2.22 µm, and measuring the mean value at 1.95–2.05 µm. That the two dither positions are different by 3 suggests there are unaccounted-for systematic uncertainties in these estimates. Given the high-quality of the data and the water ice models of S. Protopapa et al. (2014), the appearance of a strong 3-µm band, but weak or absent 1.5- and 2.0-µm bands suggests micrometer or smaller ice grains. The apparent absence of the crystalline water ice Fresnel peak at 3.1 µm could suggest the presence of amorphous water ice, but the shape and strength of this feature depends on temperature (R. M. Mastrapa et al., 2009), and likely grain size and interaction with dust. Models of 120 K micrometer-sized amorphous and crystalline ice grains show little difference around the 3.1 µm region (S. Protopapa et al., 2014). Radiative transfer modeling is needed to further test the nature of the water ice grains in 3I/ATLAS.
4 Spectral Modeling
To derive gas production rates () and rotational temperatures (), the background-subtracted IFU data for CO2, CO, and H2O were subject to spectral modeling using optimal estimation routines as part of the Planetary Spectrum Generator (PSG; G. L. Villanueva et al. 2018, including synthetic fluorescence models described by G. Villanueva et al. 2025). Initially, we constructed spectral models by taking the average spectrum for each gas inside a circular aperture centered on the nucleus. The relatively high signal-to-noise ratio (SNR) in this IFU region assisted in helping define the choice of continuum shape and the gases to be included in the model. Figure 4 shows the integrated spectra within this aperture, along with the best-fitting PSG model gas emission components. For the m region, we also identified an emission band at 4.85 m consistent with OCS, which was included in the fit.
A zeroth-order polynomial fit was sufficient for describing the continua underneath the CO2 and CO features, whereas for H2O, we experimented with various analytic functions to produce a good fit, including the wing of the 3 m ice band. Although a 5th order polynomial was found to be sufficient to reproduce the shape of the continuum in this region, a lack of formal constraints across the 2.6–2.8 m region spanning the H2O 2.7 m band led to increased uncertainties on . Therefore, for H2O, we adopted a more physically constrained continuum model, formed as the product of a linear slope and a sigmoid function:
where is the linear continuum normalization point, is the linear slope, is the y-intercept point, is the depth of the 3-µm band, controls the slope of the band edge, and controls the wavelength of the band edge. The (variable) exponential onset of the sigmoid function effectively matches the quasi-Gaussian shape of the blue wing of the 3 m amorphous H2O ice band (e.g. G. Leto & G. A. Baratta, 2003).
The CO2 and CO spectral regions were modeled using the methods described by C. E. Woodward et al. (2025). The gas outflow velocity was set at 0.44 km s-1, based on the standard relationship (e.g. T. Ootsubo et al., 2012). Pixels close to the nucleus can be affected by significant line opacity and PSF-related flux losses, which are difficult to accurately model. Therefore, we subsequently proceeded with a “-curve” analysis, deriving the production rates and rotational temperatures as a function of distance from the nucleus within successive (independent) partial annular sectors centered on the brightest (nucleus-containing) pixel. To avoid flux losses from pixels at the very edge of the IFU, and to focus on a uniform coma angular region, a inscribed angle was used for all partial annuli, with radial bounds parallel to the north-east and south-east edges of the IFU; see inset diagram in Figure 5. We found a annulus width to provide sufficient signal-to-noise for all species. Given the very high SNR for CO2, we also generated a higher-resolution curve for this species using a annulus width. The resulting production rates and rotational temperatures as a function of are shown in Figure 5.
For the m region, H2O spectral models were generated using PSG, while the continuum was determined using Equation 1, combined with the =5000 Solar spectral model of R. C. Bohlin et al. (2014). Additional (weak) spectral contributions from coma CO2 to this region were added based on our best fit to the 4.3 m band. The H2O production rate and rotational temperature were retrieved using a Gaussian processes approach within the George software package (S. Ambikasaran et al., 2015), and uncertainties were derived using a Markov chain Monte Carlo algorithm (emcee; D. Foreman-Mackey et al., 2013).
For the nucleus-centered extract (-diameter circle), our best-fitting models give s-1, s-1 and s-1 , and s-1. From the -curve analysis, “terminal” gas production rates were derived for the annular sectors furthest from the nucleus. These provide an improved view of the coma mixing ratios, avoiding the optical depth effects that impact CO2 in the nucleus-centered extract, although they could contain additional contributions from extended/distributed coma sources. The resulting (whole-coma equivalent) terminal gas production rates are s-1, s-1, s-1, and s-1.
The curves for CO2 and CO (Figure 5) level off towards larger nucleocentric distances, indicating that gas production for these species is confined within km of the nucleus. On the other hand, the H2O curve shows no clear asymptote, and the error bars allow for the possibility of continued H2O production towards the edge of the NIRSpec IFU. Therefore, we cannot rule out a contribution to H2O gas in the outer coma from sublimating icy grains, which may be expected based on our detection of coma H2O ice. A strong icy-grain source of H2O between –10,000 km from the nucleus, would be required to reconcile our terminal value with the ultraviolet OH imaging measurements of Z. Xing et al. (2025), which indicated s-1 within a 20,000 km diameter aperture.
Although 13CO2 is securely detected in the nucleus-centered extract, optical depth effects preclude the derivation of a reliable 12CO2 production rate in this region. We attempted to retrieve (13CO2) in the annular sector between – from the nucleus, but only an upper limit s-1 could be obtained, resulting in a () lower limit on the C/13C ratio of , which is formally consistent with the terrestrial value of 89 (T. B. Coplen et al., 2002). See Appendix C for further details.
5 Discussion
Our JWST NIRSpec observations reveal that 3I/ATLAS containins a substantial volatile ice inventory, with a gas coma unusually rich in CO2 relative to H2O and CO. In the absence of clear detections of other gases, it is reasonable to infer that CO2 outgassing provides the dominant driving force for 3I’s nucleus activity, and is responsible for launching dust grains away from the nucleus to produce the distinctive scattered light coma observed at m and shorter wavelengths.
For previously-observed comets in our Solar System, the relative coma abundances of CO2, CO, and H2O are known to vary widely (up to several orders of magnitude) between different comets, with some of the variability attributed to the differing relative volatilities of these species as a function of temperature (T. Ootsubo et al., 2012; M. F. A’Hearn et al., 2012; O. Harrington Pinto et al., 2022). Indeed, variation in the coma CO2/H2O and CO/H2O mixing ratios as a function of heliocentric distance is both theoretically predicted and observed (U. Marboeuf & B. Schmitt, 2014; O. Harrington Pinto et al., 2022). However, compared with previous comets observed at similar heliocentric distances (–4 au), the CO2-dominated outgassing in 3I/ATLAS appears unusual. In Figure 6, we plot the CO2/H2O coma mixing ratio measured in previous comets as a function of , and draw a log-linear trend line, fitted to the combined dataset of long- and short-period comets (excluding the peculiar outlier C/2016 R2 and the lower limit for C/2024 E1). The fit is weighted by the data uncertainties; where no uncertainties were available, an error bar of 10% was assumed.
The interstellar object 3I/ATLAS has a coma CO2/H2O ratio of , which is 16 times larger than expected for its heliocentric distance, based on the fit to previously observed cometary data (Figure 6). This corresponds to away from the trend line, and shows that 3I’s coma CO2/H2O ratio is unusually high. The only other comet known to have a CO2/H2O ratio so far outside the normal Solar System trend is C/2016 R2 (PanSTARRS) (A. J. McKay et al., 2019). C/2016 R2 is considered to be one of the most peculiar comets ever observed, as a result of its large hypervolatile content (N. Biver et al., 2018; M. A. Cordiner et al., 2022) and correspondingly high CO/H2O ratio. The CO/H2O ratio of in 3I/ATLAS, on the other hand, is more compatible with previous cometary observations, which have values between –4 au (O. Harrington Pinto et al., 2022). Intriguingly, our CO/H2O ratio is within the range of values (1.3–1.6) measured in 2I/Borisov at au, although the H2O production rate was observed to be falling rapidly around the time of those observations (D. Bodewits et al., 2020).
Due to the presence of water ice in cometary comae at 3 au (e.g. E. Lellouch et al., 1998; H. Kawakita et al., 2004; S. Protopapa et al., 2018), previous measurements of the gas CO2/H2O ratio using larger spectroscopic apertures (e.g. T. Ootsubo et al., 2012) could have been impacted by icy grain sublimation, thus reducing the observed ratio. It is difficult to assess the full impact this would have on the points around –4 au in Figure 6. We therefore recommend further observations of the volatile content of 3I/ATLAS during its post-perihelion observing window near au, to facilitate the study of its bulk ice composition.
We used the model of J. J. Cowan & M. F. A’Hearn (1979) to calculate the CO2, H2O, and CO active surface fractions, assuming an infrared emissivity of 0.95, albedo of 4% and nucleus radius of 2.8 km (D. Jewitt et al., 2025b). The Small-Bodies-Node/ice-sublimation code (M. Van Selous & M. Kelley, 2021) was used to calculate the average sublimation rate per unit area, , at = 3.32 au. The active area is derived by dividing our terminal values by , and the nucleus active fractional area is found by dividing the active area by the nucleus surface area. The lower limits on active fractional area for CO2, CO, and H2O are , , and , respectively. According to the analysis of D. Jewitt et al. (2025b), our measured CO2 active area of 2.7 km2 is sufficient to drive the development of 3I’s observed dust coma, even in the presence of large ( m) dust grains.
A lower active area for H2O than CO2 could be partly explained by the higher sublimation temperature of H2O, if the internal temperature of most of the nucleus () was in the range (but still close enough to to allow some H2O sublimation), at the time of our observations. Sublimation of H2O may therefore become more fully activated as 3I/ATLAS moves closer to the Sun (T. H. Puzia et al., 2025), in which case a more accurate picture of the nucleus composition will be obtained. The very high CO2/H2O ratio observed by JWST could therefore indicate that is lower than that experienced by typical Solar System comets at a similar . This could arise as a result of a higher albedo or lower thermal conductivity of the nucleus surface layer compared with typical comets, leading to reduced heating or heat penetration. Higher albedo could be caused by a more ice-rich surface composition than normal, whereas lower thermal conductivity could arise from the presence of a volatile-depleted crust/mantle (A. Guilbert-Lepoutre et al., 2015). The latter was hypothesized for 1I/‘Oumuamua, as a result of irradiation by cosmic rays during the object’s interstellar passage (A. Fitzsimmons et al., 2018). Accounting for the lower volatility of H2O using the (J. J. Cowan & M. F. A’Hearn, 1979) ice sublimation model, we predict CO2/H2O at au, which is still an order of magnitude larger than other comets observed near au (Figure 6).
CO2/H2O ratios greater than unity have been only rarely observed in previous comets. This is likely due to a combination of factors, including (1) relatively sparse number statistics due to the difficulty of observing comets beyond au (where H2O sublimation is strongly suppressed), and (2) the difficulty of CO2 observations in the pre-JWST era, due to telluric obscuration in the 4.3 m region. Furthermore, bulk cometary CO2/H2O ice abundances are typically less than a few tens of percent (A. C. A. Boogert et al., 2015); a median coma CO2/H2O ratio of 17% was measured by T. Ootsubo et al. (2012), and a bulk, mission-integrated value of 7% was derived for comet 67P/Churyumov–Gerasimenko (M. Läuter et al., 2020). Indeed, considering the theory that a significant fraction of cometary ice originates in the interstellar medium (P. Ehrenfreund & S. B. Charnley, 2000; M. N. Drozdovskaya et al., 2016), where CO2/H2O ratios are –50 % (A. C. A. Boogert et al., 2015; Z. L. Smith et al., 2025), a bulk CO2/H2O ratio in excess of unity for 3I/ATLAS would be surprising, perhaps hinting at an unusual, carbon-rich chemical composition for this object.
CO2 is thought to form during the interstellar and protoplanetary disk phases of star formation as a result of reactions between CO and OH on dust grain surfaces. The formation of CO2 competes with the formation of H2O from OH + H reactions (J. A. Noble et al., 2011). Under dark, non-irradiated conditions, the barrierless hydrogenation reaction resulting in H2O dominates. However, upon exposure to UV radiation and cosmic rays, physicochemical models show that CO2 may start to dominate the ice inventory (M. N. Drozdovskaya et al., 2016). S. Notsu et al. (2021) determined that CO2 ice abundances are maximized at moderate ( erg s-1) X-ray luminosities in protostellar envelopes. Various combinations of physicochemical evolutionary scenarios can theoretically produce zones with CO2/H2O in protoplanetary disks, for example: (1) beyond 30 au in the midplanes of larger, UV-irradiated, infall-dominated disks (M. N. Drozdovskaya et al., 2016), (2) at a few au over longer timescales due to cosmic ray effects (C. Eistrup et al., 2018), or (3) in elevated disk layers due to UV irradiation from the central protostar (K. Furuya et al., 2022). Furthermore, as shown by D. J. Stevenson & J. I. Lunine (1988), diffusion of sublimated gas outward across the ice line, where it subsequently freezes out, can result in significantly enhanced abundances of that ice. E. M. Price et al. (2021) modeled this effect for CO in the presence of inward-drifting icy pebbles, to explain the high CO/H2O ratio 2I/Borisov, so a similar enrichment of solid CO2 may be expected in protoplanetary disks, just outside the CO2 ice line. An intrinsically CO2-rich composition for a fraction of the interstellar object population, formed in such regions, may therefore result. Additional theoretical modeling will be required to determine whether a high CO2/H2O ratio could be compatible with an origin for 3I/ATLAS in the low metallicity, thick-disk Galactic stellar population, as suggested by M. J. Hopkins et al. (2025).
6 Conclusion
We performed JWST NIRSpec imaging spectroscopy of the interstellar object 3I/ATLAS at au on the inbound leg of its flight through the Solar System. Rovibrational emission bands were detected of CO2, CO and H2O, in addition to dust and ice solid-state features, demonstrating the presence of a substantial, gas- and ice-rich coma comparable to that of comets from our own Solar System. The CO2 band at 4.3 m was particularly strong. Gas production rates were derived using PSG spectral models, resulting in a CO2/H2O mixing ratio of , which is above the trend as a function of observed in long-period and Jupiter-family comets (excluding the peculiar C/2016 R2), and suggests the possibility of an intrinsically CO2-rich nucleus. Such a high CO2/H2O ratio has never before been observed in a comet between –4 au. The combined capabilities of the JWST and Vera C. Rubin Legacy Survey of Space and Time (Ž. Ivezić et al., 2019) will facilitate additional observations of Solar System comets at such distances, to help improve the statistics and confirm whether 3I/ATLAS is as unusual as it appears.
A low coma H2O abundance could also be implied by our data, possibly arising as a result of reduced heat penetration through an unusually thick, insulating crust or mantle. In that case, the sublimation of the less volatile H2O ice could be inhibited relative to the (more volatile) CO2 and CO ices. Further observations at distances au will be needed, to facilitate measurement of the bulk nucleus composition of 3I/ATLAS as it passes closer to the Sun and the sublimation of H2O (and other low-volatility ices) becomes more fully activated.
Appendix A Radially-Enhanced Flux Maps
Figure 7 shows “ enhanced” versions of the flux maps in Figure 2, where the dilution of the observed column density due to quasi-spherical expansion of the coma (which is proportional to , where is the sky-projected distance from the nucleus), has been divided out.
Appendix B Retrieved Gas Rotational Temperature Curves
Rotational temperatures as a function of nucleocentric distance, retrieved from independent spectral line modeling of each observed species (Section 4) are shown in Figure 8.
Appendix C 13CO2 in 3I/ATLAS
Examination of Figure 1 indicates the presence of 13CO2 alongside 12CO2, affording the first opportunity to test the 12CO2/13CO2 ratio in an interstellar object. Compared to the much stronger 12CO2 emission, the 13CO2 band is only clearly detected in the the nucleus-centered aperture. Unfortunately, the 12CO2 in this region suffers from optical depth effects, and the low spectral resolution introduces additional difficulties in disentangling the contributions from each isotopologue. We therefore focused on analyzing the spectra from the first annular sector, where optical depth effects are reduced. We used the PSG to retrieve production rates s-1 and s-1. This corresponds to a 12C/13C lower limit of 77 (3), which is formally consistent with the terrestrial value of 89. Additional observations of 13CO2 in 3I/ATLAS at higher spectral resolution and signal-to-noise will be invaluable in separating it from its optically thick 12CO2 counterpart, in order to place improved constraints on the CO2 isotopic ratio.